OpenStudy (anonymous):

For the function f(x) = x + 9, find (f o f-1)(5)

OpenStudy (anonymous):

A. 4 B. 10 C. 0 D. 6

OpenStudy (anonymous):

@IMStuck @sweetburger @Lena772

OpenStudy (imstuck):

The first thing you have to do is find the inverse function of f(x). To find the inverse function, you swap the x and the y and then solve for y. If f(x) = x + 9, then y = x + 9. The inverse of this is x = y + 9. Then solving for y gives you x - 9 = y. Put 5 in for x here and what do you get?

OpenStudy (anonymous):

5-9=-4

OpenStudy (anonymous):

@IMStuck

OpenStudy (imstuck):

I am getting that the answer is 5 and it is not one of your choices. Let me look into this for a minute, ok?

OpenStudy (imstuck):

I don't now what to do with this. if we put x through the function f, then put it through the inverse of f, we end up with x again. In this case, x is 5. So if we put 5 through the inverse of f, and then put that result through f, we come back to 5. Which is what I keep getting. Unless I am misunderstanding what you are posting...

OpenStudy (imstuck):

Maybe someone else has a different view of what your question is

geerky42 (geerky42):

Can you recheck the question? Maybe options are from wrong question? Because inverse functions are defined as \[f(f^{-1}(x)) = f^{-1}(f(x)) = x\] Composition of function and its inverse is always x. Therefore your answer should be 5.